OpenStudy (anonymous):

A railroad tunnel is shaped like a semiellipse as shown below. The height of the tunnel at the center is 58 ft and the vertical clearance must be 29 ft at a point 21 ft from the center. Find an equation for the ellipse.

OpenStudy (anonymous):

@zpupster

OpenStudy (anonymous):

@kirbykirby

OpenStudy (anonymous):

@satellite73

OpenStudy (zpupster):

b = 58, a is unknown, and the point (21,29) is on it let the equation be equation of ellipse: x^2/a^2 + y^2/b^2= 1 x^2/a^2 + y^2/58^2 = 1 21^2/a^2 + 29^2/58^2 = 1 441/a^2 = 1 - 841/3364 441/a^2 = 3/4 3a^2 = 1764 a^2 = 588 x^2/588 + y^2/3364 = 1